# Online Variational Sequential Monte Carlo

### v1

@@ 4. Theoretical results
In this section we study the limiting behavior of OVSMC and discuss its connection with batch VSMC. As explained in Section 3, the purpose of the double optimization steps in Algorithm 2 is to improve the performance of the algorithm in practical use; however, here, for simplicity, we move the parameters $\lambda$ into $\theta$, resulting in the latter also containing algorithmic parameters. Therefore, the algorithm that we theoretically analyze corresponds to Lines 814, with reparameterization function $f_{\theta_{t}}$ and $\omega_{t+1}^{i}$ depending only on $\theta_{t}$. Furthermore, for technical reasons related to the ergodicity of an extended Markov chain that we will define below and the Lipschitz continuity (in $\theta$) of its transition kernel, the analyzed algorithm repeats twice the generation of the auxiliary variable on Line 10, where one variable is used to estimate the gradient and the other to particle propagation. All details are provided in Section A in the supplement, where the modified procedure is displayed in Algorithm A.1 and we also provide all proofs.
@@
Let Assumptions A.14.2 hold. Then for every $\theta\in\Theta$, the canonical Markov chain $(Z_{t}^{\theta})_{t\in\mathbb{N}}$ induced by $T_{\theta}$ is uniformly ergodic and admits a stationary distribution $\tau_{\theta}$.
@@
For every $t\in\mathbb{N}_{>0}$, $\gamma_{t+1}^{\theta}\leq\gamma_{t}^{\theta}$. In addition, there exist constants $a>0$, $a^{\prime}>0$ and $c\in(0,1)$ such that for every $t$,
(10) $\gamma_{t}^{\theta}\leq a\gamma_{t+1}^{\theta},\quad\gamma_{t}^{\theta}-\gamma_{t+1}^{\theta}\leq a^{\prime}(\gamma_{t+1}^{\theta})^{2},\quad\gamma_{1}\leq c.$
@@
Let Assumptions A.1, 4.2, 4.4, 4.5 and 4.6 hold. Then there exist constants $b>0$ and $b^{\prime}>0$, possibly depending on $V$, such that for every $t\in\mathbb{N}$,
(11) $\mathbb{E}[\lVert h(\theta_{\tau})\rVert^{2}]\leq\frac{b+b^{\prime}\sum_{s=0}^{t}(\gamma_{s+1}^{\theta})^{2}}{\sum_{s=0}^{t}\gamma_{s+1}^{\theta}},$
@@ 5. Experiments
In this section we illustrate numerically the performance of OVSMC. We illustrate the method’s capability of learning online good proposals and model parameters, but also show that it can serve as a strong competitor to VSMC in batch scenarios. Stochastic gradients are passed to the ADAM optimizer (Kingma and Ba,, 2015) in Tensorflow 2. All the experiments are run on an Apple MacBook Pro M1 2020, memory 8GB.
@@ 5.1. Linear Gaussian SSM
Our first example is a standard linear Gaussian SSM with $\mathsf{X}=\mathbb{R}^{d_{x}}$ and $\mathsf{Y}=\mathbb{R}^{d_{y}}$ for some $(d_{x},d_{y})\in\mathbb{N}_{>0}^{2}$. More precisely, we let $m_{\theta}(\cdot\mid x)$ and $g_{\theta}(\cdot\mid x)$ be $N_{d_{x}}(Ax,S_{u}^{\intercal}S_{u})$ and $N_{d_{y}}(Bx,S_{v}^{\intercal}S_{v})$, respectively, where $A$ and $S_{u}$ are $d_{x}\times~d_{x}$ matrices, $B\in\mathbb{R}^{d_{y}\times d_{x}}$ and $S_{v}\in\mathbb{R}^{d_{y}\times d_{y}}$. We start with the simplest case $d_{x}=d_{y}=1$ and generate data under $A=0.8$, $B=1$ and $S_{u}=0.5$. We consider two cases for $S_{v}\in\{0.2,1.2\}$ corresponding to informative and more non-informative observations, respectively. The state process is initialized with its stationary distribution $N(0,S_{u}^{2}/(1-A^{2}))$. Our aim is to estimate $A$ and $S_{u}$ and, in parallel, optimizing the particle proposal. For this purpose, we let $r_{\lambda}(\cdot\mid x_{t},y_{t+1})$ be $N(\mu_{\lambda}(x_{t},y_{t+1}),\sigma_{\lambda}^{2}(x_{t},y_{t+1}))$, where $\mu_{\lambda}$ and $\sigma_{\lambda}^{2}$ are two distinct neural networks with one dense hidden layer having three and two nodes, respectively, and relu activation functions. In Figure 1 we observe the convergence of the unknown parameters for some given starting values, noticing that more informative observations yields, as expected, faster convergence. Even though values close to the true parameters are obtained already after about $10000$ iterations, the algorithm is kept running until $t=50000$ for the purpose of learning the proposal. Figure 2 shows how the effective sample size (ESS) of the particle cloud improves while $r_{\lambda}$ is being learned, to finally reach the performance of the optimal proposal. This is evident for the smallest $S_{v}$; on the other hand, when $S_{v}$ is larger, the particles are well propagated into regions of non-negligible likelihood even with the bootstrap proposal, resulting in the normalized ESS being close to one regardless. This is noticeable in Figure 3, which display conditional proposal densities for some given $(x_{t},y_{t+1})$; we see that for both choices of $S_{v}$, the learned proposal is generally close to the locally optimal one, except in the presence of unlikely outliers in the data, in which case the learned proposal tend to move, as is desirable, its mode towards that of the prior (bootstrap) kernel.
Figure 1. Parameter learning curves for the one-dimensional linear Gaussian SSM in Section 5.1, obtained using algorithm 2 with $L=5$ and $N=10000$ for $S_{v}=0.2$ (left) and $S_{v}=1.2$ (right).
@@
Next, we show that our method can also be usefully applied in the batch-data mode. Following Naesseth et al., (2018) and Zhao et al., (2022), let $d_{x}=d_{y}=10$ and let the matrix $A$ have elements $A_{ij}=0.42^{\lvert i-j\rvert+1}$. Furthermore, let $S_{u}=I$ and $S_{v}=0.5I$. We consider two parameterizations of $B$, one sparse, $B=I$, and one dense where its elements are random, independent and standard normally distributed. With these parameterizations, a data record $y_{0:T}$, $T=100$, was generated starting from an $N(0,I)$-distributed initial state. In this part, we view the model parameters $\theta$ as being known and focus on learning of the proposal. Instead of learning, as done by Naesseth et al., (2018) and Zhao et al., (2022), different local proposals for each time step, we learn a global Gaussian proposal where the mean vector and diagonal covariance matrix are functions of $x_{t}$ and $y_{t+1}$, as modeled by neural networks. These neural networks, which are time invariant, have each a single hidden layer with $16$ nodes each and the relu activation function. In this setting, we (1) processed the given observation batch $M$ times using standard VSMC, with gradient-based parameter updates between every sweep of the data, and (2) compared the result with the output of OVSMC when executed on a data sequence of length $(T+1)M$ formed by repeating the given data record $M$ times. The number of particles was equal to $L=5$ for both methods. Figure 4 displays the resulting ELBO evolutions for the two methods and some different ADAM learning rates, and it is clear from this plot that OVSMC, with its appealing convergence properties, is indeed a challenger of VSMC in this batch context.
Figure 4. ELBO evolutions of VSMC and OVSMC (each running with $L=5$) for the multivariate linear Gaussian SSM in Section 5.1 (left: $B$ sparse; right: $B$ dense).
@@ 5.2. Stochastic Volatility and RML
In this example we focus on parameter estimation and compare OVSMC with PaRIS-based RML Olsson and Westerborn Alenlöv, (2020). We consider a univariate stochastic volatility model (Hull and White,, 1987), where for $x\in\mathsf{X}=\mathbb{R}$, $m_{\theta}(\cdot\mid x)$ and $g_{\theta}(\cdot\mid x)$ are $N(\alpha x,\sigma^{2})$ and $N(0,\beta^{2}\exp(x))$, respectively, with $\alpha\in(0,1)$, $\sigma>0$ and $\beta>0$. The state process is initialized according to its stationary distribution $N(0,\sigma^{2}/(1-\alpha^{2}))$. Figure 5 reports the learning curves of the estimated parameters $\theta=(\alpha,\sigma,\beta)$ obtained with OVSMC and particle-based RML, both starting from $\theta=(0.01,0.5,0.1)$, for a data sequence generated under $\theta=(0.975,0.165,0.641)$. In both cases, the learning rate was $5\cdot 10^{-4}$. We see that OVSMC converges faster, although it fluctuates more around the true values, while the learning curve of RML is (as expected, as the RML procedure is based on particle smoothing) less noisy. Furthermore, OVSMC is not only learning $\theta$, but also optimizes online the proposal over the same family of deep Gaussian proposals as in Section 5.1. As the number of particles grows, the computational speed is also in favor of OVSMC, since the PaRIS-based RML has a computational bottleneck stemming from its inherent backward-sampling mechanism; with our implementation, OVSMC was three times faster than PaRIS-based RML in this specific case.
Figure 5. Parameter learning curves obtained with OVSMC (with $L=5$ and $N=1000$) and PaRIS-based RML (Olsson and Westerborn Alenlöv,, 2020) (with $N=1000$), for the stochastic volatility SSM in Section 5.2. In both cases the learning rate was $5\cdot 10^{-4}$.
@@ 5.3. Deep Generative Model of Moving Agent
In this section we study the applicability of OVSMC to more complex and high-dimensional models. Inspired by Le et al., (2018, Section 5.3), we produced a long and partially observable video sequence, with frames represented by $32\times 32$ arrays, showing a moving agent. The motion is similar as the one described in the referenced paper, with the difference that the agent exhibits a stationary behavior by bouncing against the edges of the image. The agent is occluded from the image whenever it moves into the central region of the frame, covered by a $16\times 30$ horizontal rectangle. The data generation is described in more detail in Section B.1 in the supplement.
As in (Le et al.,, 2018, Section C.1), we model the generative process and the proposal through the framework of variational recurrent neural networks (Chung et al.,, 2015), with a similar architecture described in more detail in Section B.2 of the supplement.
In this setting, we generated a single video sequence with $T=10^{5}$ frames, and processed the same by iterating Algorithm 2, with $N=20$ and $L=5$, $T$ times. We remark that since we have to run the algorithm for a large number of iterations, it is clearly infeasible—and possibly not even useful—to input the whole history of frames and latent states to the recurrent neural network of the model; instead, we include only a fixed window (of length $40$) of recent frames and states and discard all the previous information. This allows the average time and memory requirement per iteration to be kept at a constant value. Figure 6 displays the moving average of the incremental ELBOs $(\log(\Omega_{t}/N))_{t\in\mathbb{N}}$, where we have chosen a moving-average window of size $500$ to partially smooth both the random fluctuations and the periodical patterns of the video, and we see how this quantity increases with the number of iterations. In addition, we provide the same type of plot for two additional videos, generated with different seeds, where no learning is performed but only the incremental ELBOs are evaluated with the previously learned model. In addition, a visual illustration of the method’s capability to generate likely videos is presented in Section B.3 in the supplement.
Figure 6. The deep generative moving-agent model in Section 5.3. Moving averages with window size 500 of $\log(\sum_{i=1}^{L}\omega_{t}^{i}(\lambda,\theta)/L)_{t=0}^{T}$ (proposal learning) and $\log(\sum_{i=1}^{N}\omega_{t}^{i}(\lambda,\theta)/N)_{t=0}^{T}$ (model learning), $T=10^{5}$, produced using OVSMC with $L=5$ and $N=20$. The same quantities are shown for two test videos, with $N=20$ and $N=500$, generated by the learned model (and no learning is performed).
@@ 6. Conclusions
We have presented the OVSMC algorithm, a procedure that extends the batch VSMC to the context of streaming data. The proposed methodology allows us to learn, on-the-fly, unknown model parameters and optimal particle proposals in both standard SSM as well as and more complex generative models. Under strong mixing assumptions, which are standard in the literature, we provide theoretical support showing that the stochastic approximation scheme of OVSMC solves the same problem as VSMC for an infinite batch size. The interesting question whether the estimator produced by OVSMC is, in the case of correctly specified SSM, despite the biased approximation of the ELBO gradient, still consistent for the true parameter is left for future work.
@@
  • Intermediate results (Section A.2). In this part, we redefine the extended state-dependent Markov chain $(Z_{t})_{t\in\mathbb{N}}$ (discussed in Section 4 in the main paper), including the original SSM as well as the random variables generated by Algorithm 3. Under strong mixing assumptions, we establish, in Proposition A.4 (corresponding to Proposition 4.3 in the main paper) uniform ergodicity of the Markov transition kernel $T_{\theta}$ governing $(Z_{t})_{t\in\mathbb{N}}$ and the existence of a stationary distribution. After this, we consider the Robbins–Monroe scheme with state-dependent Markov noise targeting the mean field $h(\theta)$ defined (as in Section 4 in the main paper) as the expectation of the noisy measurement function $H_{\theta}$ under the stationary distribution of $T_{\theta}$, and prove that it is bounded. Under the assumption that the state-process and emission transition densities as well as their gradients (with respect to $\theta$) are bounded and Lipschitz continuous in $\theta$, Proposition A.6 then establishes the existence of a solution to the Poisson equation associated with the Markov kernel $T_{\theta}$.
  • Main proofs (Section A.3). Next, we introduce further assumptions regarding the Lyapunov function associated with the mean field and the step-size sequence of the stochastic update, under which we restate and prove Theorem 4.7 and Corollary 4.8 in the main paper.

### v3

@@ 4. Theoretical Results
In this section we study the limiting behavior of OVSMC and discuss its connection with batch VSMC. As explained in Section 3, the purpose of the double optimization steps in Algorithm 2 is to improve the performance of the algorithm in practical use; however, here, for simplicity, we move the parameters $\lambda$ into $\theta$, resulting in the latter also containing algorithmic parameters. Therefore, the algorithm that we theoretically analyze corresponds to Lines 8–14, with reparameterization function $f_{\theta_{t}}$ and $\omega_{t+1}^{i}$ depending only on $\theta_{t}$. Furthermore, for technical reasons related to the ergodicity of an extended Markov chain that we will define below and the Lipschitz continuity (in $\theta$) of its transition kernel, the analyzed algorithm repeats twice the generation of the auxiliary variable on Line 10, where one variable is used to estimate the gradient and the other to particle propagation. All details are provided in Appendix A, where the modified procedure is displayed in Algorithm 3 and we also provide all proofs.
@@
Let Assumptions 4.14.2 hold. Then for every $\theta\in\Theta$, the canonical Markov chain $(Z_{t}^{\theta})_{t\in\mathbb{N}}$ induced by $T_{\theta}$ is uniformly ergodic and admits a stationary distribution $\tau_{\theta}$.
@@
For every $t\in\mathbb{N}_{>0}$, $\gamma_{t+1}^{\theta}\leq\gamma_{t}^{\theta}$. In addition, there exist constants $a>0$, $a^{\prime}>0$ and $c\in(0,1)$ such that for every $t$, $\gamma_{t}^{\theta}\leq a\gamma_{t+1}^{\theta}$, $\gamma_{t}^{\theta}-\gamma_{t+1}^{\theta}\leq a^{\prime}(\gamma_{t+1}^{\theta})^{2}$ and $\gamma_{1}\leq c$.
@@
Let Assumptions 4.1, 4.2, 4.4, 4.5 and 4.6 hold. Then there exist constants $b>0$ and $b^{\prime}>0$, possibly depending on $V$, such that for every $t\in\mathbb{N}$,
$\mathbb{E}[\lVert h(\theta_{\tau})\rVert^{2}]\leq\frac{b+b^{\prime}\sum_{s=0}^{t}(\gamma_{s+1}^{\theta})^{2}}{\sum_{s=0}^{t}\gamma_{s+1}^{\theta}},$ (9)
@@ 5. Experiments
In this section we illustrate numerically the performance of OVSMC. We illustrate the method’s capability of learning online good amortized proposals and model parameters and more complex generative models, but also show that it can serve as a strong competitor to VSMC in batch scenarios. Stochastic gradients are passed to the ADAM optimizer (Kingma and Ba,, 2015) in Tensorflow 2The Python code may be found at https://bitbucket.org/amastrot/ovsmc.. All the experiments are run on an Apple MacBook Pro M1 2020, memory 8GB.
@@ 5.1. Linear Gaussian SSM
[…]
Figure 1. Parameter learning curves for the one-dimensional linear Gaussian SSM in Section 5.1, obtained using algorithm 2 with $L=5$ and $N=10000$ for $S_{v}=0.2$ (left) and $S_{v}=1.2$ (right). The means and the quantiles are calculated on the basis of 100 learning curves, each starting with a different initial value and based on independently generated observation data.
We observe an improvement of the effective sample size (ESS) of the particle cloud as $r_{\lambda}$ is converging towards the optimal proposal, as shown in Figure 8 in Appendix B. Figure 2 displays conditional proposal densities for some given $(x_{t},y_{t+1})$ in a single run; we see that for both choices of $S_{v}$, the learned proposal is generally close to the locally optimal one, except in the presence of unlikely outliers in the data, in which case the learned proposal tend to move, as is desirable, its mode towards that of the prior (bootstrap) kernel.
@@
Next, we show that our method can also be usefully applied in the batch-data mode. Following Naesseth et al., (2018) and Zhao et al., (2022), let $d_{x}=d_{y}=10$ and let the matrix $A$ have elements $A_{ij}=0.42^{\lvert i-j\rvert+1}$. Furthermore, let $S_{u}=I$ and $S_{v}=0.5I$. We consider two parameterizations of $B$, one sparse, $B=I$, and one dense where its elements are random, independent and standard normally distributed. With these parameterizations, a data record $y_{0:T}$, $T=100$, was generated starting from an $N(0,I)$-distributed initial state. In this part, we view the model parameters $\theta$ as being known and focus on learning of the proposal. Instead of learning, as done by Naesseth et al., (2018) and Zhao et al., (2022), different local proposals for each time step, we learn, in both batch and online mode, an amortized Gaussian proposal where the mean vector and diagonal covariance matrix are functions of $x_{t}$ and $y_{t+1}$, modeled by neural networks. These neural networks, which are time invariant, have each a single hidden layer with $16$ nodes each and the relu activation function. In this setting, we (1) processed the given observation batch $M$ times using standard VSMC, with gradient-based parameter updates between every sweep of the data, and (2) compared the result with the output of OVSMC when executed on a data sequence of length $(T+1)M$ formed by repeating the given data record $M$ times. The number of particles was equal to $L=5$ for both methods. Figure 3 displays the resulting ELBO evolutions for the two methods and some different ADAM learning rates, and it is clear from this plot that OVSMC, with its appealing convergence properties and reduced noise, is indeed a challenger of VSMC in this batch context.
Figure 3. ELBO evolutions of VSMC and OVSMC (each running with $L=5$) for the multivariate linear Gaussian SSM in Section 5.1 (top: $B$ sparse; bottom: $B$ dense). In each plot, which corrsponds to a particular learning rate, five independent runs och each algorithm are displayed on top of each other and compared with the target log-likelihood.
Comparison to Campbell et al., (2021)
As anticipated in Section 1, the online variational filtering (OVF) approach of Campbell et al., (2021) performs online parameter learning, although considering a variational family that is not particle-based and thus not directly comparable to the VSMC one. Even if it is algorithmically and computationally complex and includes a large number of hyperparameters to be tuned, OVF is a relevant recent work in the context of online variational learning. Thus, we used it as a benchmark for OVSMC, and let the latter solve the same problem as in Figure 1b in Campbell et al., (2021, Section 5.1). The result is displayed in Figure 4, where we demonstrate the comparability between the two methods for learning $A$ and $B$; however, we note a significant difference in the computational time, since using the code provided by Campbell et al., (2021) OVF takes about 17 hours while OVSMC uses less than one hour for a single run. Even though we are aware that part of this difference may be due to the particular implementation used, the complexity gap between the two methods is definitely non-negligible.
Figure 4. Mean absolute errors estimating $A$ (left) and $B$ (right) of five independent runs of OVF, along with the distribution of 40 independent runs of OVSMC, with $L=5$ and $N=10^{4}$, proposal kernel as described in Section 5.1, with 64 nodes in the hidden layer of each neural network, and ADAM learning rates $10^{-3}$. Here $d_{x}=d_{y}=10$, $S_{u}=0.1I$, and $S_{v}=0.25I$ and matrices $A$ and $B$ are diagonal with i.i.d. $\operatorname{Unif}(0.5,1)$-distributed elements.
@@ 5.2. Stochastic Volatility and RML
In this example we focus on parameter estimation and compare OVSMC with PaRIS-based RML (Olsson and Westerborn Alenlöv,, 2020). We consider a univariate stochastic volatility model (Hull and White,, 1987), where for $x\in\mathsf{X}=\mathbb{R}$, $m_{\theta}(\cdot\mid x)$ and $g_{\theta}(\cdot\mid x)$ are $N(\alpha x,\sigma^{2})$ and $N(0,\beta^{2}\exp(x))$, respectively, with $\alpha\in(0,1)$, $\sigma>0$ and $\beta>0$. The state process is initialized according to its $N(0,\sigma^{2}/(1-\alpha^{2}))$ stationary distribution. Figure 5 reports the learning curves of the estimated parameters $\theta=(\alpha,\sigma,\beta)$ obtained with OVSMC and particle-based RML, starting from different parameter vectors, for 27 independent data sequences generated under $\theta=(0.975,0.165,0.641)$. In both algorithms, the learning rate was $10^{-3}$.
Figure 5. Parameter learning curves obtained with OVSMC (with $L=5$ and $N=1000$) and PaRIS-based RML (Olsson and Westerborn Alenlöv,, 2020) (with $N=1000$), for the stochastic volatility SSM in Section 5.2, with learning rate $10^{-3}$.
The learning curves show similar behavior, with OVSMC converging faster than RML in several cases when estimating $\beta$, but fluctuating somewhat more when learning $\alpha$ and $\sigma$. This is to be expected, since the RML procedure is based on particle smoothing. However, we should remember that OVSMC does not use any backward sampling and in addition to learning $\theta$ also has the ability to optimize online the proposal kernel over the same family of deep Gaussian proposals as in Section 5.1. As the number of particles grows, the computational speed is also in favor of OVSMC, since the PaRIS-based RML has a computational bottleneck stemming from its inherent backward-sampling mechanism; with our implementation, OVSMC was three times faster than PaRIS-based RML in this specific case, and this ratio grows even more in favor of OVSMC as $N$ increases.
@@ 5.3. Deep Generative Model of Moving Agent
In this section we study the applicability of OVSMC to more complex and high-dimensional models. Inspired by Le et al., (2018, Section 5.3), we produced a long and partially observable video sequence, with frames represented by $32\times 32$ arrays, showing a moving agent. The motion is similar as the one described in the referenced paper, with the difference that the agent exhibits a stationary behavior by bouncing against the edges of the image. The agent is occluded from the image whenever it moves into the central region of the frame, covered by a $16\times 30$ horizontal rectangle. The data generation is described in more detail in Appendix C.1. Following Le et al., (2018, Section C.1), we model the generative process and the proposal through the framework of variational recurrent neural networks (Chung et al.,, 2015), with a similar architecture described in more detail in Appendix C.2.
In this setting, we generated a single video sequence with $T=10^{5}$ frames, and processed the same by iterating Algorithm 2, with $N=20$ and $L=5$, $T$ times. We remark that since we have to run the algorithm for a large number of iterations, it is clearly infeasible—and possibly not even useful—to input the whole history of frames and latent states to the recurrent neural network of the model; instead, we include only a fixed window (of length $40$) of recent frames and states and discard all the previous information. This allows the average time and memory requirement per iteration to be kept at a constant value.
Figures 67 show a visual illustration of the method’s ability to generate probable videos. The goal is to analyze the quality of the learned generative model by visualizing some independently sampled videos and comparing them to a reference video generated from a different seed than the one used for training. We consider a 150 frames long reference video. For each sample we use the proposal distribution to generate the latent states for the first 60 iterations, i.e., we access the true frame to sample and then decode the states, and after $t=60$ we continue with the generative model only. The images show every third frame for the true video and five independent sample videos. We note that the first 60 frames of the sample videos look similar to the true video, as expected since the model has access to the true observations. After that, the sample videos continue similarly in an initial phase, but start to move at different speeds when the agent disappears behind the rectangle after $t=75$. As the generated frame sequences show, the model predicts a smooth movement pattern of the agent, with some random perturbations, as in the second sample, where the agent leaves the rectangle after $t=96$ from the top instead of below.
Figure 6. See caption in Figure 7.
Figure 7. Comparison between original video and five samples under the learned model. Until $t=60$, frames are generated using the learned proposal and the given data to encode the latent states; after this, the frames are produced using the generative model.
@@ 6. Conclusions
We have presented the OVSMC algorithm, a procedure that extends the batch VSMC to the context of streaming data. The proposed methodology allows us to learn, on-the-fly, unknown model parameters and optimal particle proposals in both standard SSM as well as and more complex generative models. Under strong mixing assumptions, which are standard in the literature, we provide theoretical support showing that the stochastic approximation scheme of OVSMC solves the same problem as VSMC for an infinite batch size. The interesting question of whether the estimates produced by OVSMC are still consistent for the true parameter in the case of correctly specified SSM despite the biased approximation of the ELBO gradient remains open for future work. Moreover, from an application point of view, it would be appealing to explore, empirically and theoretically, the case where the parameters vary through time (i.e. when dealing with non-stationary data), with the aim of identifying the conditions under which OVSMC would still work.
@@
  • Intermediate results (Section A.2). In this part, we redefine the extended state-dependent Markov chain $(Z_{t})_{t\in\mathbb{N}}$ (discussed in Section 4), including the original SSM as well as the random variables generated by Algorithm 3. Under strong mixing assumptions, we establish, in Proposition A.4 (corresponding to Proposition 4.3) uniform ergodicity of the Markov transition kernel $T_{\theta}$ governing $(Z_{t})_{t\in\mathbb{N}}$ and the existence of a stationary distribution.
    After this, we consider the Robbins–Monroe scheme with state-dependent Markov noise targeting the mean field $h(\theta)$ defined (as in Section 4) as the expectation of the noisy measurement function $H_{\theta}$ under the stationary distribution of $T_{\theta}$, and prove that it is bounded. Under the assumption that the state-process and emission transition densities as well as their gradients (with respect to $\theta$) are bounded and Lipschitz continuous in $\theta$, Proposition A.6 then establishes the existence of a solution to the Poisson equation associated with the Markov kernel $T_{\theta}$.
  • Main proofs (Section A.3). Next, we introduce further assumptions regarding the Lyapunov function associated with the mean field and the step-size sequence of the stochastic update, under which we restate and prove Theorem 4.7 and Corollary 4.8.